Problem 3: Show U = Y-EIYIX) satisfies ElUX] = 0 and ElU2IX) = Var(Y|X)

please help me! Thanks in advance
Answer #1

3)

Now, E(U|X)=E[Y-E(Y|X)|X]=E[Y|X]-E[Y|X]=0

Also, E(U^{2}|X)=E[(Y-E(Y|X))^{2}|X]=E[(Y^{2}+E(Y|X)^{2}-2YE(Y|X))|X]

=E[Y^{2}|X]+E[(Y|X)^{2}|X]-2E(Y|X).E(Y|X)

=E[Y^{2}|X]+E[(Y|X)^{2}]-2E(Y|X)^{2}=E[Y^{2}|X]-E[(Y|X)^{2}]=Var(Y|X)

Hence Proved.